Parth (parthkohli):

If \(x^a = y^b = z^c\) and \(y^2 = zx\), then the value of \(\dfrac{1}{a} + \dfrac{1}{c}\) is?

Parth (parthkohli):

\[\dfrac{1}{a} + \dfrac{1}{c} = \dfrac{a + c}{ac}\]Any substitution?

OpenStudy (shubhamsrg):

x^2a = z^b . x^b = z^2c x^(2a-b) = z^b and z^(2c-b) = x^b or (2a-b) logx = b logz and b logx = (2c-b)log z => (2a-b)/b = b/(2c-b)

Parth (parthkohli):

The choices are: 1. \(\dfrac{2}{b}\) and, 2. \(2a\)

Parth (parthkohli):

I have eliminated two more because they definitely are not the answers.

Parth (parthkohli):

Ohhh!!

OpenStudy (shubhamsrg):

(2a-b)(2c-b) = b^2 4ac - 2ab - 2bc =0

Parth (parthkohli):

Yeah, it's \(2/b\) by hit-and-trial

OpenStudy (shubhamsrg):

yep, just divide both sides by 2abc

Parth (parthkohli):

HOW DID THAT NOT COME TO MY MIND?!

OpenStudy (shubhamsrg):

you must be busy with some of your research :)

Parth (parthkohli):

-_-

OpenStudy (shubhamsrg):



Parth (parthkohli):